### v6

# Clarify Confused Nodes via Separated Learning

Abstract
Graph neural networks (GNNs) have achieved remarkable advances in graph-oriented tasks. However, real-world graphs invariably contain a certain proportion of heterophilous nodes, challenging the homophily assumption of traditional GNNs and hindering their performance. Most existing studies continue to design generic models with shared weights between heterophilous and homophilous nodes. Despite the incorporation of high-order messages or multi-channel architectures, these efforts often fall short. A minority of studies attempt to train different node groups separately but suffer from inappropriate separation metrics and low efficiency. In this paper, we first propose a new metric, termed Neighborhood Confusion (NC), to facilitate a more reliable separation of nodes. We observe that node groups with different levels of NC values exhibit certain differences in intra-group accuracy and visualized embeddings. These pave the way for Neighborhood Confusion-guided Graph Convolutional Network (NCGCN), in which nodes are grouped by their NC values and accept intra-group weight sharing and message passing. Extensive experiments on both homophilous and heterophilous benchmarks demonstrate that our framework can effectively separate nodes and yield significant performance improvement compared to the latest methods. The source code is available in https://github.com/GISec-Team/NCGNN.
Index Terms:
Graph Neural Networks, Homophily, Heterophily, Node Classification, Semi-supervised LearningI Introduction
Graph structured data can effectively model real-world interactive systems and has received considerable attention recently. Graph neural networks (GNNs) have been proven to be powerful in many graph-related applications, such as anomaly detection [1, 2, 3], fraud detection [4, 5, 6] and recommendation systems [7, 8, 9, 10]. Traditional GNNs [11, 12, 13] are adept at learning node representations in graphs where the vast majority of nodes are homophilous, i.e., nodes are primarily connected to nodes with same labels. However, real-world graphs often contain heterophilous nodes [14] and are known as heterophilous graphs when the proportion of these nodes is high, which challenges the homophily assumption of traditional GNNs and resulting in diminished performance [15, 16, 17]. In heterophilous graphs, the patterns of connections between nodes are usually complex and varied, rather than being limited to connections with similarly labeled nodes. This is referred to as the heterophily problem, and recent studies [18, 19] have highlighted that the key to addressing this problem is to focus on the diversity of nodes in the graph. Inspired by this, it’s natural to designing flexible model architectures for characterizing diverse nodes, whether homophilous or heterophilous.
Most recent studies endeavor to design general models that still enable weight sharing between homophilous and heterophilous nodes. However, even with the incorporation of higher-order neighbors [20, 21, 22, 23] or multi-channel architectures [24, 25], these models fail to perform consistently well across all evaluation datasets. This may be due to two reasons: 1) these methods do not fundamentally alter the message-passing mechanism that essentially conforms to homophily assumption, thereby introducing noise [26]; 2) nodes within the same graph may exhibit significant behavioral differences due to distribution bias [27, 28]. For example, categorizing an interdisciplinary paper relies on more specific criteria compared to a paper confined to a single discipline [29]. Therefore, sharing weights universally across all nodes in a real-world graph is impractical, prompting the necessity for a metric to separate nodes with different characteristics. While numerous studies have explored the separation of neighbors [30, 31] or spectral signals [24, 25], the separation of target nodes or transformation weights, which we refer to as separated learning, remains under-explored. Only a minority of studies [14, 30] have focused on this separating scheme, but they suffer from incorrect separation metrics and low efficiency.
Beginning with the classical Node Homophily (NH) metric [32] derived directly from the definition of homophily or heterophily, we analyze the bipartite graph scenario as depicted in Fig. 1. According to the node homophily metric defined in Eq. (1), the homophily of any node in the bipartite graph is zero. Nevertheless, graph convolutional network (GCN) can still perform well in such extremely heterophilous scenario, presenting a paradox [33]. This phenomenon has been widely observed in recent studies [33, 34, 24], yet they neither ease this anomaly directly with their models nor explore it on a wider range of real-world graphs. Moreover, analytical papers [35, 36, 37] reach the consensus that message passing with too many layers leads to over-smoothing or indistinguishability on node representations, highlighting the importance of distinguishability over consistency with central nodes for model performance. Therefore, we propose a new metric termed Neighborhood Confusion (NC) and detail its superiority over NH. Our proposed framework is guided by this metric for node separation, further validating the effectiveness of NC. Our contributions can be summarized as follows:
We propose a new metric NC based on observations of results on GCN from multiple perspectives. And we prove that it is an appropriate metric for node separation and can be computed in real time during model training due to its low computational complexity.
We design a novel framework named Neighborhood Confusion-guided Graph Convolutional Network (NCGCN), which is the first to break the weight sharing assumption of GCN. In our framework, the metric estimation and model inference are jointly optimized in an end-to-end manner. Furthermore, we demonstrate the potential of our framework to serve as a general enhancement plug-in by substituting its backbone with an alternative message passing neural network.
Our framework achieves SOTA in ten widely used benchmarks, surpassing the latest methods tailor-made for these datasets while maintaining the simplicity of GCN. Extensive experiments demonstrate that our proposed metric NC plays a key role during model training.
II Related Work
II-A Heterophily Problem
To address the heterophily problem, several studies resort to the spectral theory. GPRGNN [20] adaptively optimizes the Generalized PageRank weights to control the contribution of propagation at each layer. Mixhop [38], H2GCN [34] and Snowball [39] combine ego- and neighbor-embedding or multi-order intermediate representations, while FAGCN [25] and ACM-GCN [24] utilize multi-channel local filters to capture local information. Limited by the globality of spectral theory, these methods are inherently unable to separate nodes and filter neighbors. Additionally, the correlation between high frequency and heterophily lacks reliability.
Spatial methods allow for fine-grained adaptations of message passing and transformation rules [40]. Geom-GCN [32] and WRGAT [41] rewire graphs to aggregate neighbors with structural similarities in latent space. SNGNN [42] and GGCN [31] construct a similarity matrix, drawing upon embeddings from the preceding layer, to separate the dissimilar neighbors. These methods require computing new topologies or similarities and thus are time and memory consuming.
Several studies also introduce new metrics, e.g., aggregated similarity score [24], local assortativity [41], 2NCS [43], density-aware homophily [19] and neighbor homophily [44]. However, these are often not effectively leveraged to guide model training, owing to high computational complexity or incompatibility with message passing layers. It reduces both the rationality of proposed metrics and the theoretical support of proposed models. To the best of our knowledge, we are the first to incorporate the proposed metric into model training in an end-to-end manner.
II-B Related Work on Label Reuse
Label distribution is the direct reason why the graphs are categorized as homophilous or heterophilous, so the label reuse [45] technique, i.e., utilizing information from labels, pseudo labels or soft assignment prediction, can be helpful. GBKSage [30] groups neighbors based on whether they belong to the same class as the target nodes, soft-separating neighbors through a gating mechanism. GHRN [46] calculates similarity of post-aggregation pseudo labels and then filters connections. CPGNN [47] and HOG [26] propagate pseudo labels to achieve an affinity matrix for calibrating the training process. Both of them use label propagation that still conforms to the homophily assumption. LWGCN [48] separates each class of neighbors and concatenates them. This approach can introduce considerable nulls, especially when a node has fewer neighbors but the graph has many classes. Furthermore, the majority of the aforementioned methods compromise the simplicity inherent in label information, introducing complexity by computing feature similarities. In contrast, our NC metric demonstrates that label distribution alone is sufficient to effectively guide models without being overly simplistic or sensitive.
II-C Related Work on Separated Learning
Separated learning involves applying distinct learning rules to target nodes. UDGNN [14] uses continual learning to achieve this separation. However, the metric NH they use is inappropriate, as we mentioned above. One possible reason for the gap between low NC and high NC groups is that they are sampled from different distributions. GraphDE [28] tries to identify in-distribution instances and treats synthesized out-of-distribution instances as noise, stopping gradient descent for the latter. In contrast, assuming that out-of-distribution instances belong to the other distribution, our approach learns the separated part with the other learner.
III Preliminaries
III-A Notations
An attributed graph can be represented as , where and are the sets of nodes and edges respectively, is the node feature matrix, and is the node label matrix. Here we use , , to denote the number of nodes, the dimension of the node features, and the number of node classes, respectively. Each row of (i.e., ) represents the feature vector of node , and each row of (i.e., ) represents the one-hot label of node . The structure elements can also be represented as an adjacency matrix that encodes pairwise connections between the nodes, whose entry if there exists an edge between and , and otherwise. Based on the adjacency matrix, we can define the degree distribution of as a diagonal degree matrix with entries representing the degree value of . We denote the set of neighbors of node within hops as .
III-B Node Homophily
The most widely used metric is Node Homophily (NH) [32]. For a target node , its NH value is defined as:
where represents the node label. NH represents the proportion of the direct neighbors that have the same label as the target node in the neighborhood. According to whether the NH is greater or less than 0.5, nodes are categorized as homophilous and heterophilous [14]. The average NH computed from the nodes across the entire graph is usually regarded as a metric for assessing graph homophily.
III-C Graph Convolutional Network
The standard GCN with two layers proposed in [11] can be formulated as:
where is the soft assignment of label predictions, and is the symmetric normalization operation with . It can be regarded as two layers of message passing, represented by , followed by transformation . Note that the message passing is indiscriminate to all neighbors and the transformation shares weight among all nodes.
IV Metrics and Observations
Heterophily is harmful to traditional GNNs, as different classes of nodes will be inappropriately mixed during message aggregation, leading to the indistinguishability of node representations [34]. However, GCN still performs well on bipartite graphs, which are completely heterophilous under the definition of NH, bringing paradox [33]. As shown in Fig. 1, means that NH metric considers the nodes in the bipartite graph to be extremely heterophilous. To alleviate this paradox, some new metrics have been proposed, such as aggregated similarity score [24], which is calculated by the similarity of intermediate representations, but with high computational complexity.
IV-A Neighborhood Confusion
Rethinking the aforementioned issue, we contend that the perceived homophily or heterophily is merely a superficial phenomenon. The crux influencing model performance lies in the difficulty of distinguishing nodes, intrinsically tied to the label distribution surrounding target nodes, as well as the non-independent nature of message passing. Therefore, we propose a new metric named Neighborhood Confusion (NC), which can measure the label diversity in the node neighborhood.
Definition 1.
Neighborhood Confusion. For a target node , its Neighborhood Confusion is defined as:
where is the set of label classes, normalizes NC to , and higher value indicates a greater degree of label confusion within target neighborhood. The receptive field of NC can be 1- or 2-hop by setting or , which is adaptively configured across different datasets.
Proposition 1.
NC metric can outline the joint label distribution in the -hop ego-net of node , and can be approximated as a loose form of joint label distribution entropy in highly heterophilous scenarios, but with reduced computational complexity by focusing on the most frequent label in the ego-net.
Proof.
The joint probability that a target node and its neighbors in the -hop neighborhood have label can be denoted as , where is subject to the label distribution of neighbors. We will abbreviate the joint probability as for simplicity. Then the joint entropy of the label distribution around (and include) target node is:
Since NC is based on the most frequent label, we express it as a loose form of joint entropy. By defining and , we have:
Thus, as the label distribution in the ego-net becomes more uniform and diverse (i.e., the environment becomes more heterophilous), NC can be interpreted as an approximation of joint entropy . The range of is determined to be based on the Maximum Entropy Principle. Consequently, NC can be normalized within the interval by dividing . While joint entropy uses the full label distribution, NC only considers the most frequent label, reducing the complexity of the computation. ∎
IV-B Observations
Rethinking the scenario of bipartite graph, as depicted in Fig. 1, suggests that our metric perceives the label distribution of node neighborhoods to have a low level of confusion, which is quite different from the result of NH. Then, we will further explain why NC surpasses NH in separating nodes from two observations. As an illustrative analysis, we train a standard GCN (2 layers with 64 hidden units, 0.5 dropout rate, and ReLU) on the public benchmark Coauthor CS [49], with the 2-hop NC and a threshold of 0.4 to separate high-NC nodes from low-NC nodes.
Observation 1.
We first calculate the NH values for nodes in the validation set and divide all nodes into ten groups according to the deciles of their NH values. Fig. 2 reports the average test accuracy of each group on 10 runs, along with the proportion of high-NC and low-NC nodes in each group. Consistent with previous studies, GCN achieves exceptionally high test accuracy within the group with the lowest NH, indicating that nodes in this group are not the least distinguishable. In conjunction with our NC metrics, this phenomenon can be easily explained by the negative correlation between test accuracy and the proportion of high-NC nodes. Specifically, the proportion of high-NC nodes in the group with NH of 0.0 is lower than that in the group with NH of 0.1, thereby resulting in a higher test accuracy. When the NH is 0.1 or higher, as it increases, the test accuracy rises as the proportion of high-NC nodes decreases. Such observation suggests that the NC metric can characterize the distinguishability of nodes to a certain extent. The results on more benchmark datasets are presented in Fig. 6.
Proposition 2.
NC metric not only measures the distinguishability of nodes in a graph but also provides insight into the classification uncertainty for each node. Nodes with higher NC values are generally more challenging to classify due to the heterophily of their neighborhoods. The classification error rate for such nodes can be theoretically bounded using NC.
Proof.
To understand how neighborhood confusion impacts node classification, we first turn to conditional entropy , which represents the expected uncertainty in the target node’s label given the labels of its neighbors . In other works, reflects how much confusion remains about the target node’s label even after knowing its neighborhood’s labels. The conditional entropy can be formulated as follows:
where represents the conditional probability that node has label given its neighborhood labels .
Assuming that the label probability of the target node is determined by its neighborhood label distribution, we have:
For the most frequent label in the neighborhood, we have:
Let . The remaining labels () occupy the proportion , and for simplicity, we assume they are uniformly distributed, such that each of these labels has a conditional probability as follows:
Substituting the above conditional probabilities into the definition of conditional entropy, we obtain:
When the neighborhood size is sufficiently large, we can approximate . This allows us to establish the connection between conditional entropy and NC.
According to the definition of NC, we have:
Then the conditional probability of the most frequent label is:
Here we consider the base of the logarithm to be 2. Then, substiituting Eq. (12) into Eq. (10) yields:
Viewing conditional entropy as a function of NC, it is easy to prove that conditional entropy is positively correlated with NC in the interval (refer to Appendix -A for more details). In other words, as the neighborhood becomes more confused (with a higher NC value), the conditional entropy increases, indicating greater uncertainty in predicting the label of the target node based on neighborhood information. Therefore, the NC metric can measure the classification difficulty (or distinguishability) of nodes in a graph.
Next, we apply Fano’s Inequality [50] to establish a theoretical lower bound on the classification error rate based on the conditional entropy as follows:
We can see that the lower bound function on the right side remains positively correlated with NC, i.e., an increase in NC will lead to a larger lower bound on the classification error rate, indicating that nodes with higher NC values experience greater classification uncertainty and pose challenges for accurate classification. Furthermore, the classification error rate can be theoretically bounded using NC. The curve of conditional entropy and classification error bound versus NC under different numbers of label categories is shown in Appendix -A. ∎
Observation 2.
Furthermore, we visualize the embedding obtained from the output of the first layer of the trained GCN in Fig. 3(a). The clusters corresponding to different classes can be roughly separated by decision boundaries depicted by black dashed lines. These black triangles symbolize high-NC nodes, predominantly distributed proximate to the cluster boundaries. When different decision boundaries (black dashed lines) intersect, clusters comprising nodes with high-NC values emerge. The clusters circled in Fig. 3(b) encompass diverse categories of high-NC nodes, thereby necessitating more complex decision boundaries for effective segregation. Test results show that the overall test accuracy is 93.99%, with 95.91% for the low-NC group and only 57.86% for the high-NC group.
These phenomena support the argument that high-NC nodes confuse GCN. We describe this as nodes becoming confused when their surrounding neighbors hold diverse views. In such cases, nodes require clarification, that is, they are highly motivated to reconsider their raw features, discern messages from their noisy neighbors, and ultimately adopt targeted classification rules. Given the significant gap between high- and low-NC nodes, it is natural to optimize them separately to fit different learning processes. With our NCGCN model for separating the two groups of nodes during learning, high-NC nodes are distanced from the decision boundaries and are more evenly distributed within the clusters of their respective classes, as shown in Fig. 3(c). The test results show that the overall test accuracy improves to 95.67%, with 95.90% for the low-NC group and 82.47% for the high-NC group, yielding a significant boost in the latter.
Proposition 3.
Node Confusion (NC) metric exhibits greater robustness compared to the Node Homophily (NH) metric. Unlike NH, which focuses more on the label consistency between the target node and its neighborhood, NC primarily considers the number of high-frequency labels. Assuming a sufficiently large number of neighbors, NC tends to be more stable when label perturbations occur in the neighborhood, whereas NH is more susceptible to uncontrollable effect, particularly when the label of the target node is disturbed. Moreover, compared to similarity-based metrics, NC can be computed efficiently during model training. Compared to label-based metrics, it is free from the constraint of consistency with the target nodes, thereby allowing a larger range of neighbors to be counted. Compared to difficulty metrics, NC does not require temporal tracking across multiple epochs, allowing for immediate node differentiation within a single training step. Compared to predictability metrics, NC encompasses all nodes, including those with high complexity, rather than excluding low-predictability or out-of-distribution samples, enabling a more comprehensive approach to node classification.
V Methodology
Based on the observations, we can conclude that NC metric can measure the distinguishability of nodes, thus being more suitable for node separation. Due to the huge gap between these two groups of nodes, it is presumed that they necessitate distinct learning processes tailored to their respective levels of distinguishability. In this section, we propose the Neighborhood Confusion-guided Graph Convolutional Network (NCGCN). Fig. 4 shows the overall framework.
V-A Neighborhood Confusion-guided Mask
In our framework, we separate the nodes into two groups, i.e., high- and low-NC nodes. Similar to [46, 51, 42], such absolute separation brings high computational efficiency. According to Eq. (3), the calculation of NC metrics relies on label information. However, it is not permissible to use all the ground-truth labels during training, so NC is calculated using pseudo labels, i.e., the model’s predicted labels for the nodes are used in place of the ground-truth labels for metric computation. Specifically, we initialize NC values of all nodes as 0, and then update NC values each time the validation accuracy reaches a new high. A threshold is used to separate nodes into high- and low-NC groups, thus yielding the Neighborhood Confusion-guided Masks (NC mask) by
where is the indicator function. Both and are diagonal matrices. Nodes with NC values below (or above) are separated into low-NC (or high-NC) groups.
V-B Neighborhood Confusion-guided Message Passing
We transform GCN with two separation strategies: 1) the separation of transformation weights at the first layer enables learners in each group to focus on specific optimization processes and eliminates noise from others; 2) the separation of message passing at the second layer prevents the outputs of different learners from being inappropriately mixed. Such separation designs bring two additional conveniences: 1) raw features can be incorporated with group-specific scalar weights; 2) the model complexity of different learners can be controlled by different dropout rates. These two kinds of separations can be easily achieved by deforming the GCN formula. Specifically, we formalize the NCGCN using the NC mask as follows:
where indicates the learning channels for different node groups, and are the weight matrices, is the activation function, is the dimension of hidden layers. In the first layer, left multiplication by implies performing row masking or target masking on , which maintains the rows of the corresponding groups in , and further realizes the weight separation, i.e., the weight is not shared between low-NC and high-NC nodes. In doing so, will aggregate all the first-order information and keep zero in the masked rows on each subsequent layer. In the second layer, right multiplication by implies performing column masking or source masking on , which can filter the biased information from different groups, so that the target node will only aggregate information from the same group of neighbors in the neighborhood:
The above two learning channels are used to characterize low-NC and high-NC nodes, respectively. Considering the importance of raw features, we introduce a linear transformation parameterized by to recover them. Then we utilize learnable scalar weight to combine the two outputs for each channel by:
Finally, we derive the soft assignment prediction through a MLP parameterized by as follows:
V-C Model Training
We employ the cross-entropy as the classification loss:
where means the sum of the diagonal elements of the matrix. The pseudo code implementation is shown in Algorithm 1. During model training, once the validation accuracy reaches a new high, the NC values and masks of all nodes will be recalculated and updated, and used in subsequent epochs. Unlike other metrics, the low computational complexity of NC metric allows it to be calculated and updated in real time during model training. The two processes of metric estimation and model inference will be alternately optimized to achieve better node classification.
Pubmed | Photo | Computers |
Cora Full | Actor | Penn94 |
#Classes | 3 | 8 | 10 | 15 | 5 | 70 | 5 | 2 | 5 | 5 | ||||||||
#Features | 500 | 745 | 767 | 6805 | 8415 | 8710 | 932 | 4814 | 2325 | 2089 | ||||||||
#Nodes | 19717 | 7650 | 13752 | 18333 | 34493 | 19793 | 7600 | 41554 | 864 | 2205 | ||||||||
#Edges | 44324 | 119081 | 245861 | 81894 | 247962 | 63421 | 26659 | 1362229 | 7754 | 46557 | ||||||||
NH | 0.87 | 0.86 | 0.82 | 0.86 | 0.93 | 0.69 | 0.62 | 0.56 | 0.33 | 0.27 | ||||||||
NC () | 0.14 | 0.07 | 0.08 | 0.06 | 0.05 | 0.09 | 0.32 | 0.68 | 0.51 | 0.60 | ||||||||
NC () | 0.20 | 0.18 | 0.20 | 0.13 | 0.10 | 0.16 | 0.43 | 0.90 | 0.58 | 0.72 | ||||||||
MLP | 86.86 0.51 | 91.28 0.63 | 84.35 1.02 | 95.22 0.17 | 95.63 0.24 | 60.60 0.77 | 36.82 1.19 | 75.34 0.64 | 31.99 4.11 | 44.61 4.23 | ||||||||
GCN | 88.08 0.50 | 93.94 0.47 | 89.72 0.43 | 94.24 0.27 | 96.31 0.15 | 70.42 0.36 | 33.15 1.75 | 76.55 0.57 | 42.63 3.64 | 47.83 2.90 | ||||||||
SAGE | 88.73 0.67 | 93.32 0.45 | 90.01 0.77 | 94.35 0.58 | 97.39 0.18 | 70.51 0.60 | 36.37 0.21 | OOM | 45.15 3.99 | 51.64 2.68 | ||||||||
GAT | 88.74 0.52 | 94.21 0.57 | 89.74 0.47 | 94.35 0.25 | 96.63 0.25 | 70.02 0.57 | 36.55 1.28 | 77.68 2.37 | 45.09 4.34 | 44.11 5.59 | ||||||||
H2GCN | 89.43 0.48 | 95.48 0.43 | 89.68 1.02 | 95.85 0.33 | 97.64 0.50 | 71.01 0.42 | 32.62 2.91 | 83.20 0.55 | 45.03 6.46 | 41.47 1.55 | ||||||||
GPRGNN | 90.46 0.75 | 93.97 0.67 | 87.92 1.96 | 95.74 0.31 | 97.77 0.15 | 71.37 0.48 | 39.22 1.17 | 84.55 0.54 | 45.15 4.95 | 45.85 2.33 | ||||||||
FAGCN | 89.16 0.67 | 94.89 0.48 | 87.61 1.17 | 95.94 0.56 | 97.37 0.33 | 69.10 1.17 | 40.90 1.12 | 76.97 0.69 | 45.20 4.45 | 53.28 1.28 | ||||||||
ACM-GCN | 90.40 0.29 | 95.04 0.44 | 90.31 0.39 | 96.10 0.22 | 97.09 0.20 | 72.19 0.36 | 39.09 1.42 | 82.98 0.63 | 43.74 4.03 | 39.64 7.84 | ||||||||
CPGNN | 87.54 1.72 | 90.49 2.43 | 83.50 2.23 | 95.42 0.24 | 97.61 0.48 | 68.47 1.02 | 33.52 1.71 | OOM | 39.18 1.28 | 28.32 2.11 | ||||||||
WRGAT | 89.13 0.53 | 92.23 0.64 | 86.13 0.76 | 95.08 0.28 | OOM | 68.67 0.37 | 37.91 1.63 | OOM | 42.46 3.71 | 45.47 3.56 | ||||||||
GBKSage | 85.49 0.61 | 81.61 8.05 | 81.20 5.62 | 95.22 0.28 | OOM | 69.19 0.47 | 37.27 0.98 | OOM | 45.32 4.61 | 45.41 3.19 | ||||||||
HOG | 85.25 0.77 | 90.88 0.68 | 82.07 1.31 | 95.94 0.16 | OOM | 60.96 0.86 | 38.00 1.72 | OOM | 41.87 5.94 | 39.92 2.38 | ||||||||
GloGNN | 84.87 0.61 | 92.41 0.72 | 86.05 1.71 | 91.54 5.14 | 97.54 0.53 | 67.04 0.75 | 32.88 2.03 | 83.07 1.35 | 44.39 3.99 | 49.07 8.68 | ||||||||
SNGNN | 86.85 0.62 | 90.79 2.59 | 85.30 1.02 | 95.52 0.28 | 96.43 0.10 | 71.11 0.31 | 38.68 0.78 | 73.87 1.01 | 41.17 3.61 | 38.27 2.62 | ||||||||
CAGNN | 87.54 0.67 | 92.75 1.29 | 89.59 0.63 | 94.50 0.81 | 97.93 0.30 | 69.88 0.53 | 39.83 0.85 | 79.49 2.00 | 43.86 2.98 | 53.28 2.26 | ||||||||
NCGCN | 91.64 0.53† | 95.45 0.45 | 90.81 0.46 | 96.64 0.29† | 98.63 0.24 | 73.42 0.58† | 43.16 1.32 | 84.74 0.28 | 46.84 4.36 | 54.15 1.47 | ||||||||
NCSAGE | 91.55 0.38 | 95.93 0.36† | 90.43 0.72 | 96.48 0.25 | 98.69 0.26† | 72.58 0.65 | 43.89 1.33† | 81.77 0.71 | 48.07 3.47 | 54.42 1.41 | ||||||||
[…]
w/o Mess. Separ. | 90.93 0.49 | 93.69 0.61 | 87.10 0.56 | 95.21 0.38 | 96.83 0.67 | 67.98 0.60 | 39.97 2.16 | 76.30 0.40 | 32.57 3.18 | 36.72 4.08 | ||||||||
NH Separ. | 90.88 0.56 | 94.54 0.42 | 89.71 0.53 | 95.75 0.41 | 98.06 0.40 | 66.17 0.73 | 39.03 1.63 | 78.52 0.51 | 39.24 6.64 | 31.81 5.08 | ||||||||
[…]
V-D Plug-in Property and Efficiency
While we focus on separate learning research on the classic message passing architecture of GCN, it’s notable that our NC-guided framework possesses a plug-in property, i.e., it can be combined with any other message passing neural networks by simply replacing the backbone model. To show the plug-in property of our framework, we further replaced GCN with another message passing neural network, SAGE [12], forming the NCSAGE model. Its performance will be demonstrated in the experimental section.
In addition, our framework introduces three key optimizations at design time, which include: 1) constructing a hash dictionary to efficiently index the -hop neighbors of a node, enabling the estimation of NC values to operate with linear complexity; 2) computing the two separated channels in parallel; 3) using threshold-based separated learning to make the learner of each channel sparser than the corresponding backbone in terms of parameters. These architectural features avoid imposing excessive computational overhead, and ultimately ensure good computational efficiency.
VI Evaluations
VI-A Experimental Settings
VI-A1 Datasets
We use 10 real-world benchmarks from [20, 49, 52], among which the first five (Pubmed, Photo, Computers, Coauthor CS and Coauthor Physics) are homophilous and the last five (Cora Full, Actor, Penn94, Chameleon and Squirrel) are heterophilous. Note that such division is based on the criterion of . We adopt de-duplicated version of the last two, namely Chameleon Filtered and Squirrel Filtered, as provided in [53]. For all datasets, we adopt dense data splitting as in [20], i.e., randomly split them into training, validation and testing sets with a proportion of 60%/20%/20%. We provide more details for these datasets in Appendix -B.
VI-A2 Competitors
We compare our methods with a total of 15 competitors in three categories, including classical methods such as two-layer MLP, GCN [11], SAGE [12] and GAT [13], spectral methods such as H2GCN [34], GPRGNN [20], FAGCN [25] and ACM-GCN [24], spatial methods such as CPGNN[47], WRGAT [41], GBKSage [30], HOG [26], GloGNN [54], CAGNN [44] and SNGNN [42]. The methods are the latest and tailor-made for the heterophily problem. We provide more details for these competitors in Appendix -C.
VI-A3 Setup
To ensure the stability and reproducibility of the results, we use ten consecutive seeds to fix the data splitting and model initialization. For all methods, we set the maximum epochs to 500 with 100 epochs patience for early stopping, the hidden dimension to 512, and the optimizer to Adam [55]. For common parameters like learning rate, weight decay and dropout rate, we search in the same parameter space if their codes are publicly available. The key parameters of NCGCN/NCSAGE like hop and threshold will be searched in their respective parameter spaces and . Detailed parameter settings are reported in Appendix -D.
VI-B Analysis
To clearly organize this section, we answer six Research Questions (RQ) and demonstrate our arguments by extended experiments. For simplicity, we only discuss the validity of our framework on NCGCN in RQ2 - RQ6.
VI-B1 RQ1: Does our framework outperforms baselines and its backbone across homophilous and heterophilous datasets?
Table I reports the node classification results of all methods on ten benchmarks, from which we can draw the following conclusions: 1) Among all baselines, spectral methods show a dominant performance advantage on the heterophily problem. However, although they achieve relatively high overall performance, they lack a certain degree of universality, e.g., GPRGNN does not perform well on Photo and Computer. This is mainly due to the fact that the high-order prior of these methods is not applicable to all datasets, and the weight sharing hinders further performance improvement. 2) Spatial methods usually have complex architectural designs and large hyperparameter spaces, leading to higher computational complexity as well as unstable and poor performance. For instance, WRGAT, GBKSage and HOG are out of memory (OOM) on Coauthor Physics and Penn94. 3) As a spatial method, our framework achieves the best overall performance on two types of datasets without being out of memory (OOM), and reaches first place on the public leaderboard in six well-explored benchmarks (refer to Appendix -E), indicating its performance advantage. Additionally, our framework shows an average performance improvement of 8.04% and 4.54% on their respective backbones (GCN and SAGE), confirming its general plug-in property. Notably, the sparsification of the adjacency matrix by the NC-mask allows our NCSAGE model to overcome the OOM problem encountered by SAGE on Penn94. All these phenomena highlight the effectiveness of our framework in alleviating heterophily problem, especially the superiority of our proposed NC metric and priors.
VI-B2 RQ2: How do the key designs in NCGCN contribute?
Our framework features three key designs: 1) the separation of transformation weights in the first layer; 2) the separation of message passing in the second layer; 3) the guidance of separation by the NC metric. To further investigate the effectiveness of these key designs, we conduct three ablation studies: 1) removing separation strategies (target masking and source masking) and degenerating NCGCN to an adaptive mixture of GCN output and raw features (w/o Separ.); 2) restoring inter-group message passing (source masking) in the second layer (w/o Mess. Separ.); 3) guiding separation using NH metric and a threshold of 0.5 (NH Separ.). The results of ablation studies are shown below the dashed line in Table I. By comparing w/o Separ. and w/o Mess. Separ. with NCGCN, we conclude that introducing separation strategies enable more high-NC nodes to benefit from a more appropriate feature learning process, especially in heterophilous graphs where the separation strategies yield higher gains. Additionally, NH Separ. outperforms w/o Separ. on homophilous graphs but fails on heterophilous ones, indicating that NH metric has limited generalizability on heterophilous graphs, reinforcing the superiority of our NC-guided separation strategy. In conjunction with the node distributions under different metrics in Fig. 5, it is evident that NC and NH hold different insights on the confusion of nodes in the graph. Our NC metric facilitates a more effective re-adjustment of node distribution, particularly in heterophilous graphs. In summary, the performance gap between ablation models and our NCGCN highlights the effectiveness of the three key designs.
VI-B3 RQ3: Whether the NC metric can measure the distinguishability of nodes?
Before analyzing the interpretability of NC metric, we first review the perplexing phenomenon exhibited by NH metric in Fig. 6. We can see that the grouped test accuracy curves based on NH metric consistently present a ‘ticking’ trend, especially where the node groups with extremely low homophily show high test accuracy. It cannot be explained well by the definition of NH metric. However, there seems to be a certain negative correlation between test accuracy and the proportion of high-NC nodes, offering an insight into this phenomenon’s interpretation.
We further group nodes based on NC metric and record the grouped test accuracy curves of different models, as shown in Fig. 7. Firstly, we note that in most cases, the grouped test accuracy curves monotonically decrease as NC value increases. This common trend can be reasonably explained as follows. When nodes have neighbors with multiple label types, they exhibit higher NC values according to the definition of neighborhood confusion. In such cases, nodes receive more heterophilous information during message passing, which can introduce noise, thereby reducing their distinguishability. This decreased distinguishability might present greater challenges in classifying these nodes, consequently affecting test accuracy. Therefore, we can conclude that NC metric can, to some extent, measure the distinguishability of nodes, with nodes having higher NC values being less distinguishable.
Furthermore, it is observable that compared to some baseline methods, our separation learning framework typically improves the classification performance of different NC groups more effectively. In a weight-sharing learning process, the idea of simplifying the model conflicts with the eliminating biases. This conflict causes interference between the different patterns of low-NC and high-NC nodes, forcing the model to make a trade-off between fairness and accuracy. In contrast, our separation learning framework breaks the assumption of weight sharing, offering customized learners for both low-NC and high-NC nodes, thus holding the potential for debiasing.
VI-B4 RQ4: Whether metric estimation and model inference complement each other?
Our framework relies on predicted pseudo-labels to estimate NC values, which subsequently guide the separate learning process. On the one hand, the accuracy of the pseudo-labels predicted by the model, i.e., the effectiveness of model inference, will impact the accuracy of metric estimation. On the other hand, the accuracy of metric estimation will impact the NC mask and subsequently impact node separation results, finally impacting both model training and inference. Therefore, we further analyze the interplay between metric estimation and model inference. We first initialize , meaning all nodes are initially assigned to the low-NC group, and then commence model training. Fig. 8 shows the curves of average test accuracy and the average pseudo NC value across 10 runs. Note that we use real labels to calculate the ground-truth NC value at the global level (red dashed line), and use the predicted pseudo-labels to estimate the pseudo NC value (black curve). In the early stages of training, we observe continuous fluctuations in the model’s test accuracy and the estimated NC value. With the increase of training epochs, the NC value estimated by the model gradually align more closely with the ground-truth NC value, and the model’s test accuracy progressively improves. Both aspects tend to stabilize after a certain number of training epochs, particularly the metric estimation results which, in most cases, converge to the ground-truth. This phenomenon suggests that the processes of metric estimation and model inference are synergistically optimized during training, complementing each other and enhancing the overall performance.
Additionally, it’s noteworthy that our framework is initialized with no high-NC nodes, but the estimated NC values can eventually converge to the ground truth NC, suggesting that the node separation task (i.e., metric estimation) is easier to optimize than the node classification task (i.e., model inference). This is a significant advantage of our proposed metric and framework: even when the model cannot achieve high prediction performance, our metric can still robustly separate the nodes in a reasonable manner.
VI-B5 RQ5: How do the two hyperparameters of hops and threshold impact NCGCN?
Metric computation and node separation in our framework are controlled by the two key parameters, hops and threshold . To further investigate the impact of these hyperparameters on NCGCN, we fix the other parameters to their optimal settings and plot the test accuracy curves under varying and , as shown in Fig. 9.
When we fix , we observe that NCGCN exhibits relatively small performance fluctuations with varying on the five homophilous graphs and Cora Full. This is because these datasets have relatively stable label distributions, and the NC values computed under different parameter settings are more consistent, resulting in similar node groupings. Consequently, NCGCN is more robust to changes in the hop parameter. In contrast, the neighborhood of different hops in heterophilous graphs will exhibit more diverse label distributions, resulting in significantly distinct node groupings. As a result, NCGCN demonstrates heightened sensitivity to parameter changes. For large-scale heterophilous graphs like Penn94, the 1-hop receptive field already includes a sufficient number of neighbors. The main reason for its poor performance in the 2-hop setting lies in the ineffective utilization of 2-hop information by the target nodes to evaluate their own state.
When we fix , we observe that NCGCN exhibits relatively small performance fluctuations with varying on the five homophilous graphs and Cora Full. This is because these datasets typically have a large proportion of low-NC nodes, as shown in Fig. 5, and as the threshold increases, the node groupings remain stable, leading to consistent model performance. For small-scale heterophilous graphs like Actor, Squirrel Filtered, and Chameleon Filtered, NCGCN is more sensitive to threshold changes due to the more discrete distribution of nodes across different NC values. For larger heterophilous graphs like Penn94, the distribution of nodes with different NC values tends to increase with higher NC values. As the threshold increases, the division between low-NC and high-NC nodes becomes more balanced, which gradually improves the performance of NCGCN ().
VI-B6 RQ6: Whether separate learning setting exhibit different feature fusion preferences in different node groups?
In Eq. (17), we introduce learnable scalar weight for each learner to fuse aggregated features with raw features. To further explore whether learners acting on different node groups have distinct preferences in feature fusion, we plot the kernel density estimates of for the low- and high-NC groups across ten runs, as shown in Fig. 10. It is evident that the peak of the weight distribution for the low-NC group is right-skewed compared to that of the high-NC group. This implies that the learner in the low-NC group assigns greater weight to aggregated features and less to raw features during feature fusion, while the opposite is true for the high-NC group. This phenomenon aligns with intuition and can be reasonably explained as follows. Nodes in the high-NC group have neighborhoods with more complex label distributions, and the aggregated features obtained through message passing contain more heterophilous noise. Therefore, the learner for the high-NC group tend to rely more on the nodes’ raw features for representation, resulting in relatively smaller . Conversely, nodes in the low-NC group have neighborhoods with purer label distributions, and the learner for this group effectively utilizes the features obtained through message passing for node representation, resulting in relatively larger . In summary, the gap between the weights of two groups within one graph suggests that our separation strategy is necessary and effective.
ogbn-arxiv | Flickr | |
#Remark | large, temporal | large |
#Classes | 40 | 7 |
#Features | 128 | 500 |
#Nodes | 169343 | 89250 |
#Edges | 1166243 | 334120 |
NH | 0.64 | 0.32 |
NC () | 0.04 | 0.37 |
NC () | 0.06 | 0.57 |
MLP | 55.39 0.17 | 46.91 0.09 |
GCN | 55.87 0.18 | 47.82 0.11 |
SAGE | 57.52 0.29 | 47.87 0.08 |
GAT | 56.81 0.23 | 47.28 0.16 |
H2GCN | 55.23 0.14 | 48.26 0.11 |
GPRGNN | 57.29 0.28 | 48.03 0.10 |
FAGCN | 57.03 0.27 | 48.12 0.36 |
ACM-GCN | 56.78 0.19 | 47.54 0.28 |
CPGNN | OOM | OOM |
WRGAT | OOM | OOM |
GBKSAGE | OOM | OOM |
HOG | OOM | OOM |
GloGNN | OOM | OOM |
SNGNN | 56.21 0.18 | 47.11 0.13 |
CAGNN | 50.34 3.49 | 47.01 0.10 |
NCGCN | 58.29 0.27 | 48.43 0.09 |
NCSAGE | 59.20 0.32 | 48.33 0.13 |
VI-B7 Efficiency Comparison
To investigate the computational efficiency of different methods, we conduct a relatively fair comparison of actual time consumption on Pubmed, as shown in Fig. 11. For all methods, the number of {hidden units, layers} are set to {64, 2}, respectively. The results demonstrate that our NCGCN outperforms partial spatial methods in terms of computational efficiency.
First, CPGNN, GBKSage and HOG employ two-phase training and soft rewiring techniques, resulting in significant time consumption. WRGAT constructs a weighted multi-relational graph before training, rendering it equivalent to multi-channel GATs with reduced efficiency. These spatial methods typically involve additional computational steps for adjusting message passing rules, such as rewiring, auxiliary tasks, or gating mechanisms, thereby increasing the computational burden. In contrast, the NCGCN architecture is essentially a simple deformation of GCN, while also incorporates the three key designs mentioned in Sec. V-D. Consequently, it avoids excessive computational overhead, ultimately ensuring good computational efficiency. Please refer to Appendix -F for more analysis. Additionally, although some spatial methods consume less computation time, they have certain design flaws in their model architectures. For GloGNN, directly inputting the adjacency matrix into an MLP contradicts permutation invariance. For SNGNN, the approach of selecting the most similar neighbors seems to remain confined within the homophily assumption. CAGNN combines raw and aggregated features without considering label heterogeneity. These design flaws result in these methods falling short in terms of accuracy.
VI-B8 Performance Analysis on Large-scale Graphs with Inductive Learning
In this section, we further analyze the performance of various models on large-scale graphs with inductive learning, focusing on the ogbn-arxiv [56] and Flickr [57] benchmarks. These two datasets present distinct challenges: ogbn-arxiv contains temporal information and a higher degree of homophily, while Flickr is more heterophilous, providing a diverse testing ground for models designed to address the heterophily issue in GNNs. As shown in Table II, the proposed NCGCN and NCSAGE, which are guided by the NC metric, deliver the best performance across both datasets. These results demonstrate the effectiveness of the NC metric in separating nodes based on neighborhood label diversity, allowing the models to apply distinct learning processes to nodes with different levels of neighborhood confusion. This separation is especially crucial for handling the heterophilous nature of Flickr while maintaining high performance on the larger and more homophilous ogbn-arxiv dataset. Additionally, it is important to note that several complex models, such as GBKSAGE and HOG, experience timeout and OOM errors, further demonstrating the challenge of scaling certain spatial methods to large-scale graphs. In contrast, NCGCN and NCSAGE handle these large-scale datasets efficiently, illustrating the scalability of the NC-guided separation framework in real-world graph applications.
VII Conclusions and Outlook
Based on the observations of intra-group accuracies and node embeddings, we find that nodes within a single real-world graph exhibit certain differences. It indicates that the weight-sharing scheme in GCN is not reliable. Thus we explore a spatial method with separated learning, NCGCN, which is totally guided by our proposed metric NC. Our method addresses two common problems in existing spatial GNNs based on node separation: 1) the separation metrics are not appropriate; 2) training steps and separation strategies are complex and inefficient.
Our framework is simple yet effective enough to be a new start for spatial GNNs. The backbone can be replaced by more message passing models. The robustness and debiasing potential are also open to be explored. Furthermore, NC is an easy-to-compute metric that can effectively separate nodes for different optimization processes. This metric can also be further utilized in data mining, predictability, and interpretability. The limitation of NCGCN is that it utilizes only second-order information following traditional GNNs. It seems sufficient on most real-world datasets except for long-distance dependent ones. Another problem is that high-NC nodes are relatively rare on homophilous graphs, and we can use graph data augmentation [58] for this issue in the future.